OpenStudy (anonymous):

Square root of 1 - cos20degrees / 2. I've got it simplified to the square root of Sin of 20 degrees over 2 but I don't know where to go from there

OpenStudy (jmark):

We know that Square root of ((1 - cos A)/ 2) = sin^2 (A/2) Here A = 20 So Square root of (1 - cos20degrees / 2) = Square root of (sin^2 (20degrees/2)) = Square root of (sin^2 (10degrees)) = sin (10 degrees) = 0.17